OpenStudy (anonymous):

50 grams of propane (c3h8) is oxidized (combusted/burned) completely to co2 and h20 by reaction with 300 grams of pure oxygen (o2). What is the composition of the reaction products in grams? grams of c02= grams of h20= grams of c3h8= grams of 02=

OpenStudy (rushwr):

first write the balanced equation

OpenStudy (anonymous):

i did but i am still confused like i balanced it like this: C3H8+5o2-->3Co2+4h20

OpenStudy (rushwr):

yep

OpenStudy (rushwr):

find the no. of moles of C3H8 moles = mass divided by molar mass

OpenStudy (anonymous):

what should i do after i get the moles

OpenStudy (rushwr):

moles of CO2 = 3 times of moles of C3H8

OpenStudy (rushwr):

use the same equation moles = mass/ molar mass

OpenStudy (rushwr):

then we can find mass

OpenStudy (rushwr):

of CO2

OpenStudy (anonymous):

should i find the moles of all of them first

OpenStudy (rushwr):

yeah yeah

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

now how do i find the mass

OpenStudy (rushwr):

mass = molar mass * moles

OpenStudy (anonymous):

the masses i got don't add up to 350 idk y

OpenStudy (anonymous):

thank you though :)

OpenStudy (rushwr):

wait I'll check for u

OpenStudy (rushwr):

Mass of CO2 is 150 g approximately

OpenStudy (rushwr):

Mass of H2O = 81.8 g approximately

OpenStudy (rushwr):

Looking at the stoichiometric ratio we can see the ratio of moles of C3H8: O2 is 1:5 That means 1 moles of C3H8 reacts with 5 moles of O2 We know the no. of moles of C3H8 that was 1.136 right? That means to oxidise C3H8 completely we need (1.136 *5) moles right? so no. of moles of O2 needed = 5.68 moles But we are said that we have 300 g of O2 Using that mass we should find the no. of moles of O2 . The total O2 present Total moles = 300/32 = 9.375 But to complete the reaction we only need 5.68 moles of O2 That means O2 is in excess. So that will remain in the products side. The remaining would be 9.375 - 5.68 = 3.695 moles Now we can find the mass of O2 remaining mass = 3.695 * 32= 118.24g

OpenStudy (anonymous):

oh okay thank you so much

OpenStudy (rushwr):

No problem ... hope you got it :)